if i multiply the number by 4 and then add 5, i get the same answer as when i double the number and then add 8
Answer by checkley79(3341)

4N+5=2N+8
4N-2N=8-5
2N=3
N=1.5 ANS.
PROOF:
4*1.5+5=2*1.5+8
6+5=3+8
11=11